# FedReFT: Federated Representation Fine-Tuning with All-But-Me Aggregation

### v1

@@ C.1 Commonsense Reasoning
We train and evaluate our models on eight commonsense reasoning datasets spanning different types of open-ended QA tasks, following Hu et al. (2021a), we construct all examples. Table 10 shows the dataset samples.
@@
  • BoolQ Clark et al. (2019): A yes/no question answering dataset consisting of naturally occurring questions. We remove the associated passages to ensure a fair comparison.
@@
  • PIQA Bisk et al. (2020): A dataset for physical commonsense reasoning. The model must select the more plausible solution to everyday physical tasks.
@@
  • SIQA Sap et al. (2019): Focuses on social interaction reasoning by asking the model to choose responses based on human intent and consequences.
@@
  • HellaSwag Zellers et al. (2019): Requires choosing the most coherent sentence completion given a context, often involving physical or temporal common sense.
@@
  • WinoGrande Sakaguchi et al. (2021): Inspired by the Winograd Schema Challenge Levesque et al. (2012), this dataset contains fill-in-the-blank problems with binary choices requiring commonsense coreference reasoning.
@@
We follow the experimental setup in Hu et al. (2021a) by fine-tuning our models on a combined training corpus referred to as COMMONSENSE170K, which merges all of the above datasets. Evaluation is conducted individually on each dataset’s test split.
@@
Table 10: Examples from commonsense reasoning tasks: BoolQClark et al. (2019), PIQABisk et al. (2020), HellaSwagZellers et al. (2019), and SIQASap et al. (2019). Each instruction is followed by the answer selected during evaluation.
@@
Table 11: Examples from math reasoning tasks: AQUA Ling et al. (2017), GSM8K Cobbe et al. (2021), and SVAMP Patel et al. (2021). Each instruction is followed by the correct answer derived through step-by-step reasoning.
@@
C.2 Arithmetic Reasoning
We evaluate arithmetic reasoning using seven benchmark datasets that cover a range of math word problem types. As in Hu et al. (2021a), we construct all examples without using golden or retrieved passages. Data samples are shows in Table 11.
  • AQuA Ling et al. (2017): Presents algebraic word problems in a multiple-choice format.
  • GSM8K Cobbe et al. (2021): A widely used benchmark of grade-school math problems requiring multi-step reasoning.
  • SVAMP Patel et al. (2021): A more challenging dataset that tests robustness to paraphrased and structurally altered word problems.
Following Hu et al. (2021a), we train our models on a combined training set named MATH10K.
C.3 Natural Language Understanding
For NLU, we evaluate on the GLUE benchmark following the evaluation protocol in Wu et al. (2024b). Data samples for shown in Table 12.
Table 12: Examples from GLUE benchmark  Wang et al. (2018) tasks: MNLI, SST-2, QNLI, and QQP.
@@ C.3 Natural Language Understanding
Appendix D Evaluation Loss-Based $\alpha$ Tuning
To personalize the blending of local and aggregated intervention parameters in FedReFT, we employ an evaluation loss-based tuning strategy for the mixing coefficient $\alpha\in[0,1]$. This coefficient governs how much each client integrates the aggregated All-But-Me (ABM) intervention parameters with its own local updates:
$\theta_{k}^{\text{new}}=(1-\alpha)\cdot\theta_{k}^{\text{local}}+\alpha\cdot\theta_{k}^{\text{ABM}},$ (10)
where $\theta_{k}$ can represent the LoReFT intervention components $W$, $R$, and $b$ for client $k$.
Tuning Procedure:
  1. 1.
    For a set of candidate $\alpha$ values ($\alpha\in\{0.0,0.1,\ldots,1.0\}$), the client computes interpolated parameters.
  2. 2.
    For each $\alpha$, the client evaluates its performance using a local validation set and records the evaluation loss $\mathcal{L}(\alpha)$.
  3. 3.
    The optimal mixing coefficient is selected by:
    $\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(\alpha).$ (11)
Appendix E Theoretical Foundation: Geometric Median via Weiszfeld’s Algorithm
This optimization is non-smooth and convex, and generally lacks a closed-form solution. However, Weiszfeld’s algorithm Weiszfeld (1937) provides an efficient iterative method to approximate $\mathbf{y}^{*}$. We now derive and justify this algorithm via the Majorization-Minimization (MM) framework.
We define the cost function to be minimized:
$f(\mathbf{y})=\sum_{i=1}^{n}\|\mathbf{y}-\mathbf{x}_{i}\|_{2}.$ (13)
This function is convex but non-differentiable at points where $\mathbf{y}=\mathbf{x}_{i}$. Weiszfeld’s algorithm avoids such points during updates by construction.
Solving the above yields the Weiszfeld update rule:
The update is only valid when $\mathbf{y}^{(k)}\neq\mathbf{x}_{i}$ for all $i$, a condition that can be enforced by initialization and step-size dampening if needed.
From MM theory Lange (2016), each iteration satisfies:
$f(\mathbf{y}^{(k+1)})$ $\leq Q(\mathbf{y}^{(k+1)}|\mathbf{y}^{(k)})$ (17)
$\leq Q(\mathbf{y}^{(k)}|\mathbf{y}^{(k)})=f(\mathbf{y}^{(k)}),$
E.1 Application in FedReFT: ABM Aggregation
We compute this using Weiszfeld’s algorithm for each parameter type independently, ensuring robustness to outlier clients and misaligned updates. This enables stable and personalized aggregation without sacrificing task-specific semantics.
Weiszfeld’s algorithm provides a theoretically grounded and computationally efficient way to compute the geometric median, making it ideal for ABM aggregation in heterogeneous FL. By leveraging this algorithm in FedReFT, we ensure robustness in aggregation and improve both convergence and personalization in non-i.i.d. federated environments.
Appendix F Ablation Study
F.1 Comparison of Aggregation method on different task
We use only 20% of the training data from the COMMONSENSE170K dataset, split among three clients, and evaluate the models using the SIQA task. Figure 3 shows that Geometric Median ABM aggregation outperforms all other approaches. Similarly, we split 50% of the MATH10K dataset among five clients, train each client for only five local epochs, and evaluate the results using the GSM8K evaluation set. Additionally, we use RTK GLUE for the natural language understanding (NLU) task.
Figure 3: Comparison of Aggregation Strategies Across Tasks. Performance of FedAvg, Mean-ABM, and Geometric Median-ABM on three benchmark task groups: Commonsense Reasoning, Arithmetic Reasoning, and GLUE. Geometric Median-ABM consistently outperforms FedAvg and Mean-ABM, highlighting its robustness and effectiveness in heterogeneous federated settings.
Model Total P(M) # TP(M) TP%
LLaMa-1.1B 1100.05 0.72 0.0655
LLaMA 7B 6,738.42 2.10 0.0311
LLaMA-2 7B 6,738.42 2.10 0.0311
LLaMA-3 8B 8,030.27 2.10 0.0261
LlaMa-2-13B 13,015.86 6.55 0.0503
RoBERTa Large 355.36 0.0492 0.0138
Table 14: We vary LLaMA model sizes with $C=3$ clients following the DT design for commonsense reasoning task, alongside a centralized LoReFT baseline. As model capacity increases, we observe notable performance gains, with the largest model approaching the accuracy of the centralized setting. First four experiments on the Standalone centralize setup and later four experiments on the FL setup.
Method BoolQ PIQA HellaS. Avg $\uparrow$
LLaMa 7b 69.30 84.4 93.1 82.27
LLaMa-2 7B 71.10 83.8 94.3 83.07
LLaMa-3 7B 75.1 90.2 96.3 87.20
Tiny LLaMa 1B 63.83 49.18 46.03 53.01
LLaMa 7B 65.84 77.75 67.64 70.41
LLaMa-2 7B 68.93 74.81 77.73 73.82
LLaMa-3.2 3B 64.12 77.47 81.40 74.33
LLaMa-3 7B 72.60 85.85 89.85 82.77
Appendix G Communication Efficiency
G.1 Additional Experimental Validation
In this section, we also conducted some additional experiments to show the robustness of FedReFT in different setups. Appendix Table LABEL:tbl:second_setup_commonsense depicts these.

### v2

@@ F.1 Commonsense Reasoning
We train and evaluate our models on eight commonsense reasoning datasets spanning different types of open-ended QA tasks, following Hu21, we construct all examples. Table 18 shows the dataset samples.
@@
  • BoolQ clark2019boolq: A yes/no question answering dataset consisting of naturally occurring questions. We remove the associated passages to ensure a fair comparison.
@@
  • PIQA bisk2020piqa: A dataset for physical commonsense reasoning. The model must select the more plausible solution to everyday physical tasks.
@@
  • SIQA sap2019socialiqa: Focuses on social interaction reasoning by asking the model to choose responses based on human intent and consequences.
@@
  • HellaSwag zellers2019hellaswag: Requires choosing the most coherent sentence completion given a context, often involving physical or temporal common sense.
@@
  • WinoGrande sakaguchi2021winogrande: Inspired by the Winograd Schema Challenge levesque2012winograd, this dataset contains fill-in-the-blank problems with binary choices requiring commonsense coreference reasoning.
@@
We follow the experimental setup in Hu21 by fine-tuning our models on a combined training corpus referred to as Commonsense170K, which merges all of the above datasets. Evaluation is conducted individually on each dataset’s test split.
@@
Table 18: Examples from commonsense reasoning tasks: BoolQ clark2019boolq, PIQA bisk2020piqa, HellaSwag zellers2019hellaswag, and SIQA sap2019socialiqa. Each instruction is followed by the selected answer during evaluation.
@@
Table 19: Examples from math reasoning tasks: AQuA ling2017program, GSM8K cobbe2021training, SVAMP patel2021nlp, and MAWPS koncel2016mawps. Each instruction is followed by the correct answer derived through step-by-step reasoning.
@@
MAWPS Solve the following math word problem:
Tom has 3 boxes of pencils, each containing 12 pencils. He gives 7 pencils to his friend. How many pencils does Tom have left in total?
29
Table 20: Examples from GLUE benchmark wang2018glue tasks: MNLI, SST-2, QNLI, and QQP. Each instruction is followed by the corresponding ground truth label.
@@
F.2 Arithmetic Reasoning
We evaluate arithmetic reasoning using seven benchmark datasets that cover a range of math word problem types. As in Hu21, we construct all examples without using golden or retrieved passages. Data samples are shows in Table 19.
  • AQuA ling2017program: Presents algebraic word problems in a multiple-choice format.
  • GSM8K cobbe2021training: A widely used benchmark of grade-school math problems requiring multi-step reasoning.
  • SVAMP patel2021nlp: A more challenging dataset that tests robustness to paraphrased and structurally altered word problems.
  • MAWPS koncel2016mawps: A large repository of math word problems aggregated from multiple sources, covering diverse arithmetic and algebraic reasoning tasks expressed in natural language.
Following Hu21, we train our models on a combined training set named MATH10K.
F.3 Natural Language Understanding
For NLU, we evaluate on the GLUE benchmark following the evaluation protocol in Wu2024. Data samples for shown in Table 20.
Appendix G Computational Resources
All experiments are executed on a single NVIDIA A100-SXM4-80GB GPU, except for LLaMA-2 13B, which is run on a GPUH200x8 141GB system to accommodate the computational demands of large-scale federated fine-tuning.